OpenStudy (anonymous):

let A=M2(Z) be the ring of 2×2 integral matrices the identity of A, IA

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

is the question asking for the identity matrix?

OpenStudy (anonymous):

YES

OpenStudy (misty1212):

it is the usual one

OpenStudy (anonymous):

HMM. 

OpenStudy (misty1212):

\[\begin{array}{cc} 1 & 0 \\ 0 & 1 \end{array}\]

OpenStudy (misty1212):

yeah that one

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Let G be a group and H1 , H2 normal subgroups of .one of these is a normal subgroup of G

OpenStudy (misty1212):

whew i thought it was going to be some hard ring question, not a nice easy one

OpenStudy (misty1212):

you lost me there, was it perhaps a copy and paste fail?

OpenStudy (anonymous):

1)H1 INTERCEPT H2 2)H1 UNION H2 3)H1-H2 4)AXB

OpenStudy (misty1212):

if \(A_1,H_2\) are normal in \(G\) then \(H_1\cap H_2\) is as well, it is a straightforward check

OpenStudy (misty1212):

depending of course on what your definition of a normal subgroup is there are a couple

OpenStudy (anonymous):

OK MA. HAVE SOME MORE...

OpenStudy (misty1212):

wait, you don't have to prove it, just pick one?

OpenStudy (anonymous):

If R and R'are rings, a mapping ϕ:R→R′ ring ho morphism if any of these happen ∀a,b,∈R

OpenStudy (anonymous):

ϕ(a+b)=ϕ(a)+ϕ(b) ϕ(a/b)=ϕ(a)−ϕ(b) ϕ(a.b)=ϕ(a)ϕ(b) A and C only

OpenStudy (misty1212):

wow an abstract algebra class with multiple guess questions? no proofs just pick?

OpenStudy (anonymous):

IM VERY CONFUSE and the book i have does note contain all these.

OpenStudy (misty1212):

guess, i bet you get it on the first try or google ring homorphisms one hint, no one says division is even DEFINED in a ring

OpenStudy (anonymous):

yes. i thought as much. it is multiplication and addition . right?

OpenStudy (misty1212):

or just read the top line here https://en.wikipedia.org/wiki/Ring_homomorphism

OpenStudy (misty1212):

yes A and C

OpenStudy (anonymous):

An isomorphism of a ring is both an epimorphism and ________________ Monomorphism Endomorphism Automorphism homomorphism

OpenStudy (anonymous):

i think it is homomorphism

OpenStudy (misty1212):

i am not sure what "homomorphism" of a ring means, a homo from on ring to another?

OpenStudy (misty1212):

isomorphism means a homomorphism that is both injective and surjective, or in this language "epi" and "mono"